Average Senior High School annual cost of tuition fee for all private schools last year was P46,300. A random sample of costs this year for 45 private schools indicated that the sample mean was P47,800 and a sample standard deviation was P5,600. At the 0.10 level of significance, is there sufficient evidence to conclude that the cost has increased?


1
Expert's answer
2022-05-06T13:06:03-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu\le 46300"

"H_a:\mu>46300"

This corresponds to a right-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.10," "df=n-1=44" degrees of fredom, and the critical value for a right-tailed test is "t_c =1.30109."

The rejection region for this right-tailed test is "R = \{t: t >1.30109\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{47800-46300}{5600/\sqrt{45}}\approx1.79684"

Since it is observed that "t = 1.79684> 1.30109=t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for right-tailed, "df=44" degrees of freedom, "t=1.79684" is "p = 0.039614," and since "p= 0.039614<0.10=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"

is greater than 46300, at the "\alpha = 0.10" significance level.